Two trains A and B 125m and 100 long are moving in opp directions on parallel tracks the velocity of train B is three times that of A the trains take 4 s to pass each other thenvel           velocity of each train is?

Grade:11

1 Answers

Aflah
7 years ago
let velocity of A be vvelocity of A wrt B = 4vtotal distance travelled by either trains=225mi.e velocity of A= 225/4*4 =14m/s (approx).i.e velocity of B = 42m/s